OpenStudy (anonymous):

When using nCr, how do you solve (1/3) C 1 ?

OpenStudy (dumbcow):

i dont think you can because , one factorials are only defined for positive integers, and two, 1 > 1/3 and n mustr be greater than r

OpenStudy (anonymous):

I thought there was another formula for it. Because my teacher told me (1/3)C2 was (1/3((1/3)-1)/2!)

OpenStudy (dumbcow):

oh ok, i have never seen fractions used with combinations i won't work on a calculator

OpenStudy (dumbcow):

it*

OpenStudy (zarkon):

(1/3) C 1 =1/3

OpenStudy (anonymous):

I'm pretty sure it's negative 1/3. But if it's like that, how would you do (1/2) C 2?

OpenStudy (zarkon):

no it's 1/3 but (1/2) C 2=-1/8

OpenStudy (anonymous):

Could you write up the equation for me please?

OpenStudy (zarkon):

\[\binom \alpha k = \frac{\alpha(\alpha-1)(\alpha-2)\cdots(\alpha-k+1)}{k(k-1)(k-2)\cdots 1}\]

OpenStudy (zarkon):

\[_\alpha C_k=\binom \alpha k\]

OpenStudy (anonymous):

I understand the bottom part as a factorial, but how would you do the top part?

OpenStudy (zarkon):

\[_{1/2}C_2=\frac{\frac{1}{2}\left(\frac{1}{2}-1\right)}{2(2-1)}=\frac{\frac{1}{2}\cdot\frac{-1}{2}}{2}=-\frac{1}{8}\]

OpenStudy (anonymous):

how do you how many times you repeat the "(x-1)"? You said it was α(α−1)(α−2).... but you only repeated 2 times above.

OpenStudy (zarkon):

because k=2